An electronic company in Laguna manufactures resistors that have a mean resistance of 120 ohms and a standard deviation of 13 ohms. Find the probability that a random sample of 40 resistors will have an average resistance greater than 114 ohms. 


1
Expert's answer
2022-03-18T15:06:47-0400

We have a normal distribution, "\mu=120, \sigma=13, n=40."

Let's convert it to the standard normal distribution,

"z=\cfrac{\bar{x}-\mu}{\sigma/\sqrt{n}}" =

"=\cfrac{114-120}{13/\sqrt{40}}=-2.92,"

"P(\bar{X}>114)=1-P(\bar{X}<114)=\\
=1-P(Z<-2.92)="

"=1-0.0018=0.9982" ​(from z-table)